### v1

# Entanglement-enabled advantage for learning a bosonic random displacement channel

Abstract
We show that quantum entanglement can provide an exponential advantage in learning properties of a bosonic continuous-variable (CV) system. The task we consider is estimating a probabilistic mixture of displacement operators acting on $n$ bosonic modes, called a random displacement channel. We prove that if the $n$ modes are not entangled with an ancillary quantum memory, then the channel must be sampled a number of times exponential in $n$ in order to estimate its characteristic function to reasonable precision; this lower bound on sample complexity applies even if the channel inputs and measurements performed on channel outputs are chosen adaptively. On the other hand, we present a simple entanglement-assisted scheme that only requires a number of samples independent of $n$, given a sufficient amount of squeezing. This establishes an exponential separation in sample complexity. We then analyze the effect of photon loss and show that the entanglement-assisted scheme is still significantly more efficient than any lossless entanglement-free scheme under mild experimental conditions. Our work illuminates the role of entanglement in learning continuous-variable systems and points toward experimentally feasible demonstrations of provable entanglement-enabled advantage using CV quantum platforms.
Quantum science and technology holds promise to revolutionize how we understand and interact with nature, enabling computational speedups [1], classically impossible communication tasks [2, 3], and measurements with unprecedented sensitivity [4, 5, 6]. Rapid progress during the noisy intermediate-scale quantum (NISQ) era [7] has brought these promises closer to reality, but the challenge remains to demonstrate rigorous quantum advantage for practical problems.
Over the past few years, there has been ongoing theoretical and experimental progress in exploring quantum computational advantage [8, 9, 10, 11, 12, 13, 14, 15, 16]. Another recent line of research seeks quantum advantage in learning [17, 18, 19, 20, 21, 22, 23, 24], revealing that access to quantum memory enables us to learn properties of nature more efficiently. Specifically, Refs. [18, 19] establish a framework for proving exponential separation in sample complexity between learning with and without a coherently controllable quantum memory. In contrast to its computational counterpart, this entanglement-enabled advantage in learning can be proven without invoking computational assumptions and can sometimes be more experimentally accessible. A proof-of-principle experiment has been conducted on Google’s superconducting quantum processor using $40$ qubits [18].
Most learning tasks studied so far are restricted to discrete-variable (DV) systems. It is natural to ask whether entanglement-enabled advantage can also be realized for learning properties of bosonic continuous-variable (CV) systems. This is particularly interesting and important because CV systems are ubiquitous in nature and have many applications in quantum information science, such as quantum sensing [25, 26, 27, 6, 8]. However, generalizing the results in DV systems to CV systems can be difficult because bosonic systems have infinite-dimensional Hilbert spaces, making it challenging to formulate rigorous results concerning the complexity of learning properties of these systems. Recent progress has been achieved in studies of entanglement-enhanced learning of CV-state characteristic functions [28]; however, the lower bounds obtained so far apply to a restricted class of learning strategies rather than to general entanglement-free schemes.
In this work, we rigorously establish an entanglement-enabled advantage in learning a probabilistic mixture of $n$-mode displacement operations, called a bosonic random displacement channel. Specifically, we show that any schemes without ancillary quantum memory require a number of samples exponential in $n$ to learn the characteristic functions of the channel with reasonably good precision and high success probability. On the contrary, we present a simple scheme utilizing entanglement with ancillary quantum memory (i.e., entanglement-assisted) that can complete the same learning task with a sample complexity independent of $n$, given access to two-mode squeezed vacuum (TMSV) states with sufficiently large squeezing parameter and Bell measurements (BM). This establishes an exponential separation between learning with and without entanglement in the bosonic system. The two learning scenarios are illustrated in Fig. 1. Note that our hardness results hold for arbitrarily high-energy input states and arbitrary measurements, while the presented entanglement-assisted scheme only requires a finite-energy TMSV and BM.
Furthermore, we analyze the robustness of this entanglement-enabled advantage under realistic experimental conditions. Specifically, we study the photon-loss effect, the most common noise source in optical platforms. Our results suggest that for squeezing parameters and loss rates achievable in a state-of-the-art bosonic experiment platform, the separation in sample complexity remains significant. Therefore, we anticipate that an experimental demonstration of entanglement-enabled advantage in CV quantum systems can be achieved in the near future.
Problem Setup.— We consider the task of learning an $n$-mode random displacement channel characterized by a probability distribution $p(\alpha)$ with $\alpha\in\mathbb{C}^{n}$, which transforms an input state $\hat{\rho}$ as
$\Lambda(\hat{\rho})=\int d^{2n}\alpha~{}p(\alpha)\hat{D}(\alpha)\hat{\rho}\hat{D}^{\dagger}(\alpha),$ (1)
where $\hat{D}(\alpha)\mathrel{\mathop{:}}\nobreak\mkern-1.2mu=\otimes_{i=1}^{n}\hat{D}(\alpha_{i})$ and $\hat{D}(\alpha_{i})\mathrel{\mathop{:}}\nobreak\mkern-1.2mu=\exp(\alpha_{i}\hat{a}_{i}^{\dagger}-\alpha_{i}^{*}\hat{a}_{i})$ is the displacement operator for the $i$th mode. The random displacement channel can also be equivalently described by the characteristic function of $p(\alpha)$, i.e., its Fourier transform, as (see SM S1 [29] for the derivation)
$\Lambda(\hat{\rho})=\frac{1}{\pi^{n}}\int d^{2n}\beta~{}\lambda(\beta)\operatorname{Tr}[\hat{\rho}\hat{D}(\beta)]\hat{D}^{\dagger}(\beta),$ (2)
$\lambda(\beta)\mathrel{\mathop{:}}\nobreak\mkern-1.2mu=\int d^{2n}\alpha~{}p(\alpha)e^{\alpha^{\dagger}\beta-\beta^{\dagger}\alpha}.$ (3)
Here, because of the Fourier relation, $\lambda(\beta)$ with a large $\beta$ contributes to rapidly oscillating $p(\alpha)$. Since the domain of $\beta$ is infinite in principle, we will focus on a restricted finite domain specified later. The goal is to learn the channel by estimating the characteristic function $\lambda(\beta)$. We emphasize that the goal is to characterize the channel, as opposed to identifying a particular displacement that is drawn from the distribution $p(\alpha)$. The value of $\beta$ for which $\lambda(\beta)$ is to be estimated is revealed only after all measurements are completed.
We focus on the separation between two types of learning schemes for the random displacement channel distinguished by whether or not the scheme uses entanglement between the system and an ancilla, as illustrated in Fig. 1. Throughout this work, we define an entanglement-free scheme to be both ancilla-free and concatenation-free, i.e., the output of the channel is measured destructively after each channel use. However, the entanglement-free scheme is allowed to be adaptive; for each channel use, the input to the channel and the measurement performed on the output may depend on measurement outcomes obtained in earlier rounds. This scenario is similar to Refs. [17, 22]. Several recent works have obtained lower bounds on learning DV channels that hold even with concatenation [23, 30, 31], but we will not analyze the consequences of concatenating CV channels in this work for simplicity.
Figure 1: Schemes for learning an $n$-mode random displacement channel $\Lambda$. (a) TMSV+BM, a specific entanglement-assisted (EA) scheme. (b) General entanglement-free (EF) scheme. Here we assume no concatenation is allowed, i.e., each copy of the channel acts on some input state $\rho_{0}$ and is measured destructively by some POVM $\{E\}$. The input state and measurement are allowed to be adaptively chosen depending on previous outcomes. An example of EF scheme is Vacuum+Heterodyne (see the main text).
Schemes.— Now, we present an entanglement-assisted scheme (see Fig. 1) inspired by a similar scheme to which has been proposed in DV Pauli channel estimation in Ref. [23]. Consider an $n$-mode random displacement channel $\Lambda_{B}$ acting on the $n$-mode system $B$. To learn this channel, we prepare $n$ CV Bell states with a finite squeezing parameter $r$, which is a two-mode squeezed vacuum (TMSV) state, and half of the states go through the channel while the other half stays in quantum memory. Finally, we measure the output state by CV Bell measurement (BM), which can be implemented by passing through a 50:50 beam splitter and performing homodyne measurement on output ports along different quadratures [27]. Formally, the BM POVM element labeled by $\{\zeta\in\mathbb{C}^{n}\}$ has the following form: $(I\otimes\hat{D}(\zeta))|\Psi\rangle\langle\Psi|(I\otimes\hat{D}^{\dagger}(\zeta))/\pi^{n}$; here $|\Psi\rangle$ denotes the tensor product of $n$ infinitely squeezed TMSV states, each proportional to $\sum_{k=0}^{\infty}|k\rangle|k\rangle$ when expressed in the Fock basis. To see how to learn a random displacement channel using this TMSV+BM scheme, we invoke the probability of obtaining outcome $\zeta$ from BM (see SM S2 A [29] for the derivation):
$p_{EA}(\zeta)$ $=\frac{1}{\pi^{2n}}\int d^{2n}\alpha~{}\lambda(\alpha)e^{-e^{-2r}|\alpha|^{2}}e^{\alpha^{\dagger}\zeta-\zeta^{\dagger}\alpha}.$ (4)
Fourier transforming to invert this relation, we obtain
$\lambda(\beta)$ $=e^{e^{-2r}|\beta|^{2}}\int d^{2n}\zeta~{}p_{EA}(\zeta)e^{\zeta^{\dagger}\beta-\beta^{\dagger}\zeta}$ (5)
$\mathrel{\mathop{:}}\nobreak\mkern-1.2mu=e^{e^{-2r}|\beta|^{2}}\lambda_{EA}(\beta).$
This expression indicates that, by sampling $N$ measurement outcomes $\{\zeta^{(i)}\}_{i=1}^{N}$ from a TMSV+BM scheme, one can obtain an unbiased estimator $\tilde{\lambda}(\beta)$ of $\lambda(\beta)$ by defining $\tilde{\lambda}(\beta)\mathrel{\mathop{:}}\nobreak\mkern-1.2mu=\frac{1}{N}e^{e^{-2r}|\beta|^{2}}\sum_{i=1}^{N}e^{\zeta^{(i)\dagger}\beta-\beta^{\dagger}\zeta^{(i)}}$. Note that the same set of samples can be used to estimate $\lambda(\beta)$ for different values of $\beta$ just by modifying the estimator. Using the Hoeffding’s bound, we prove the following theorem (see SM S2 A for the proof):
Theorem 1.
For any $n$-mode random displacement channel $\Lambda$, after the TMSV+BM scheme with squeezing parameter $r$ has learned from $N$ copies of $\Lambda$, and then received a query $\beta\in\mathbb{C}^{n}$, it can provide an estimator $\tilde{\lambda}(\beta)$ of $\Lambda$’s characteristic function $\lambda(\beta)$ such that $|\tilde{\lambda}(\beta)-\lambda(\beta)|\leq\epsilon$ with probability at least $1-\delta$, with the number of samples $N=8e^{2e^{-2r}|\beta|^{2}}\epsilon^{-2}\log 4\delta^{-1}$.
Let us compare the TMSV+BM scheme with a particular entanglement-free scheme that uses the vacuum state as input and heterodyne detection (Vacuum+Heterodyne). Here, heterodyne detection is defined as a projection onto the (overcomplete) basis of coherent states, i.e., $|\zeta\rangle\langle\zeta|/\pi^{n}$ with $\zeta\in\mathbb{C}^{n}$. Though it may not be the optimal entanglement-free scheme, this specific scheme helps us understand the limitations of entanglement-free schemes, which we capture more generally in Theorem 2 below.
In this scheme, the probability of obtaining POVM outcome $\zeta$ is (see SM S2 B [29])
$p_{{VH}}(\zeta)=\frac{1}{\pi^{2n}}\int d^{2n}\alpha~{}\lambda(\alpha)e^{\alpha^{\dagger}\zeta-\zeta^{\dagger}\alpha}e^{-|\alpha|^{2}}.$ (6)
In fact, the Vacuum+Heterodyne scheme can be understood as the TMSV+BM scheme with $r=0$. Inverting this relation by Fourier transforming, we may express the channel’s characteristic function in terms of the measurement probability distribution:
$\lambda(\beta)=e^{|\beta|^{2}}\int d^{2n}\zeta~{}p_{{VH}}(\zeta)e^{\zeta^{\dagger}\beta-\beta^{\dagger}\zeta}\mathrel{\mathop{:}}\nobreak\mkern-1.2mu=e^{|\beta|^{2}}\lambda_{VH}(\beta),$ (7)
which yields another unbiased estimator $\tilde{\lambda}(\beta)\mathrel{\mathop{:}}\nobreak\mkern-1.2mu=\frac{1}{N}e^{|\beta|^{2}}\sum_{i=1}^{N}e^{\zeta^{(i)\dagger}\beta-\beta^{\dagger}\zeta^{(i)}}$ given $N$ samples $\{\zeta^{(i)}\}_{i=1}^{N}$. Comparing to (5), we see that the $r$-dependent prefactor is missing from (7). Specifically, if we confine $\beta$ to $|\beta|^{2}\leq\kappa n$ with a constant $\kappa>0$, we obtain upper bounds on the sample complexity for achieving an error $\epsilon$ with success probability $1{-}\delta$ using each scheme:
$N_{EA}$ $=O(e^{2e^{-2r}\kappa n}\epsilon^{-2}\log\delta^{-1}),$ (8)
$N_{VH}$ $=O(e^{2\kappa n}\epsilon^{-2}\log\delta^{-1}).$ (9)
In particular, if we choose the squeezing parameter as $r=\Omega(\log n)$, the sample complexity $N_{EA}=O(\epsilon^{-2}\log\delta^{-1})$ of the entanglement-assisted scheme becomes independent of the number of modes $n$, while our upper bound on sample complexity of the entanglement-free scheme increases exponentially with $n$. Since the accessible squeezing parameter is bounded in practice, though, we will compare the sample complexities of the two schemes when $r$ is an $n$-independent constant below.
To illustrate the difference, we compare TMSV+BM and Vacuum+Heterodyne strategies with an example in Fig. 2. We consider a single-mode channel for ease of visualization, characterized by
$p(\alpha)$ $=\frac{2\sigma^{2}}{\pi}e^{-2\sigma^{2}|\alpha|^{2}}[\cos^{2}(\alpha_{r}\gamma_{i}-\alpha_{i}\gamma_{r})$
$~{}~{}~{}~{}~{}+\sin^{2}(\alpha_{r}\gamma_{r}+\alpha_{i}\gamma_{i})],$ (10)
$\lambda(\beta)$ $=e^{-\frac{|\beta|^{2}}{2\sigma^{2}}}+\frac{1}{4}e^{-\frac{|\beta-\gamma|^{2}}{2\sigma^{2}}}+\frac{1}{4}e^{-\frac{|\beta+\gamma|^{2}}{2\sigma^{2}}}$
$~{}~{}~{}~{}~{}-\frac{1}{4}e^{-\frac{|\beta-i\gamma|^{2}}{2\sigma^{2}}}-\frac{1}{4}e^{-\frac{|\beta+i\gamma|^{2}}{2\sigma^{2}}},$ (11)
with $\sigma=0.3$, $\gamma_{r}=1.6$, $\gamma_{i}=0$ ($\gamma\mathrel{\mathop{:}}\nobreak\mkern-1.2mu=\gamma_{r}+i\gamma_{i}$), and $r=2$ for the TMSV+BM scheme. The figure, where we present the underlying output probability distributions and their characteristic functions from Eqs. (4),(5),(6), and (7), clearly shows that in the TMSV+BM scheme with a sufficiently large squeezing parameter, the resultant probability distribution and characteristic function are almost identical to the ideal case. However, for the Vacuum+Heterodyne scheme, the vacuum noise distorts the initial probability distribution so significantly that we cannot see the signal clearly, which thus makes it harder to estimate the original characteristic function.
Figure 2: Comparison between (a) the true distribution, (b) TMSV+BM, and (c) Vacuum+Heterodyne strategies. The left panel represents the probability distribution of the true distribution and measurement probability distributions for each scheme. The right panel represents the characteristic function of probability distributions.
Lower bound.— Our upper bound on the sample complexity of the Vacuum+Heterodyne scheme scales exponentially with $n$. Can this scaling be improved using more advanced entanglement-free schemes, such as homodyne or general-dyne detection [32, 27], or by non-Gaussian resources like GKP states [33] or photon-number resolving measurements [34]? Here, using information-theoretic methods, we prove an exponential sample complexity lower bound for any entanglement-free scheme. This highlights the indispensable role of entanglement for efficiently learning bosonic random displacement channels. Our result is as follows:
Theorem 2.
Let $\Lambda$ be an arbitrary $n$-mode random displacement channel ($n\geq 8$) and consider an entanglement-free scheme that uses $N$ copies of $\Lambda$. After all measurements are completed, the scheme receives the query $\beta\in\mathbb{C}^{n}$ and returns an estimate $\tilde{\lambda}(\beta)$ of $\Lambda$’s characteristic function $\lambda(\beta)$. Suppose that, with success probability at least $2/3$, $|\tilde{\lambda}(\beta)-\lambda(\beta)|\leq\epsilon\leq 0.24$ for all $\beta$ such that $|\beta|^{2}\leq n\kappa$. Then $N\geq 0.01\epsilon^{-2}(1+1.98\kappa)^{n}$.
Here, the choice of success probability $2/3$ is arbitrary and can be easily amplified. Comparing with the entanglement-assisted sample complexity given in Eq. (8), Theorem 2 establishes a separation exponential in $n$ for cutoff coefficient $\kappa=O(1)$ and squeezing parameter $r=\Omega(\log n)$. The intuition underlying this theorem is that displacement operators $\hat{D}(\beta)$ do not generally commute with each other. Consequently, entanglement-free measurements can resolve $\lambda(\beta)$ for only a small portion of $\beta$ space. We sketch the proof below and leave the full details to SM S3 [29].
Proof Sketch.
Our proof extends the techniques of Refs. [23, 18] to the CV case. We begin by defining the following family of “$3$-peak” random displacement channels $\bm{\Lambda}^{\epsilon,\sigma}_{\mathrm{3\mhyphen peak}}=\{\Lambda_{\gamma}\}_{\gamma\in\mathbb{C}^{n}}$ whose characteristic functions and distributions of displacement are, respectively,
$\lambda_{\gamma}(\beta)$ $=e^{-\frac{|\beta|^{2}}{2\sigma^{2}}}+2i\epsilon_{0}e^{-\frac{|\beta-\gamma|^{2}}{2\sigma^{2}}}-2i\epsilon_{0}e^{-\frac{|\beta+\gamma|^{2}}{2\sigma^{2}}},$ (12)
$p_{\gamma}(\alpha)$ $\propto e^{-2\sigma^{2}|\alpha|^{2}}\left(1+4\epsilon_{0}\sin(2(\gamma_{i}\alpha_{r}-\gamma_{r}\alpha_{i}))\right).$ (13)
with positive parameters $\sigma$ and $\epsilon\mathrel{\mathop{:}}\nobreak\mkern-1.2mu=0.98\epsilon_{0}$. We will show that, even with the prior knowledge that the channel is from this family, it is still hard for entanglement-free schemes to complete the learning tasks.
The key idea is to reduce learning to binary hypothesis testing. Consider the following game between Alice and Bob: Alice chooses one of two hypotheses with equal probability: (1) Set $\Lambda=\Lambda_{0}$; (2) Set $\Lambda=\Lambda_{\gamma}$, for $\gamma$ sampled from a zero-mean homogeneous Gaussian distribution whose variance is determined by $\kappa$. Next, Alice allows Bob to use the channel $\Lambda$ $N$ times, and Bob uses his favorite entanglement-free scheme to learn from these channel uses. After Bob has finished all quantum measurements and keeps only classical data, Alice reveals some auxiliary information to Bob, who is then asked to decide whether Alice has chosen (1) or (2).
Given a learning scheme satisfying the assumptions of Theorem 2, Bob can guess correctly with high probability. This means that the outcome distributions of Bob’s scheme under hypotheses (1) and (2) must have a sufficiently large total variation distance (TVD). On the other hand, we can upper bound the contribution from each use of $\Lambda$ to the TVD to be exponentially small, where we use a technique inspired by Ref. [35] which derived the maximum fidelity of Gaussian random displacement channels. Therefore, the number of channel uses $N$ must be exponentially large to ensure a large enough TVD, which gives us the desired lower bound. ∎
Figure 3: (a) Comparison of TMSV+BM (with different loss rates), Vacuum+Heterodyne, and the entanglement-free lower bound at $\kappa=1$. The task is to estimate any $\lambda(\beta)$ such that $|\beta|^{2}\leq\kappa n$ with precision $\varepsilon=0.2$ and success probability $31-\delta=2/$ 3. The orange region represents a rigorous advantage over all entanglement-free schemes. The blue region represents an advantage over noiseless Vacuum+Heterodyne. (b) Comparison of the TMSV+BM scheme with squeezing parameter $r=1.0$ and loss rate $0.11-T=$ 0.1 with the entanglement-free lower bound of Theorem 2. (See SM S3 A for further practical considerations.) The task is the same as (a). The brown solid contour lines represent the sample complexity of TMSV+BM given by Theorem 3. The blue dashed contour lines represent the ratio of sample complexity between the entanglement-free lower bound and TMSV+BM, indicating the entanglement-enabled advantage.
Effect of loss.— Now, for practical applications, we study how the entanglement-assisted scheme is affected by photon loss, a dominant noise source in optical platforms (see SM S2 D for a discussion of more general noise models, such as phase diffusion). Photon loss transforms the relevant bosonic operator $\hat{a}$ to $\sqrt{T}\hat{a}+\sqrt{1-T}\hat{e}$, where $T$ is the transmission rate and $\hat{e}$ is the environmental mode, i.e., $1-T$ is the loss rate. We consider two different places where the loss occurs: one is before applying the channel with loss rate $1-T_{b}$ to model the preparation imperfection, and the other is after applying the channel and before the perfect BM with loss rate $1-T_{a}$, which models the finite efficiency of detection, i.e., an imperfect BM [27]. As before, we derive the relation between the measurement probability distribution and the characteristic function of the channel (with appropriate rescaling of the phase):
$\lambda(\beta)$ $=e^{e^{-2r_{\mathrm{eff}}}|\beta|^{2}}\int d^{2n}\zeta~{}p_{loss}(\zeta)e^{(\zeta^{\dagger}\beta-\beta^{\dagger}\zeta)/\sqrt{T_{a}}},$ (14)
where we define an effective squeezing parameter
$r_{\text{eff}}\mathrel{\mathop{:}}\nobreak\mkern-1.2mu=-\frac{1}{2}\log\left(T_{b}e^{-2r}+(1-T_{b})+\frac{1-T_{a}}{T_{a}}\right),$ (15)
which incorporates the loss rates. Because loss degrades the advantage from squeezing, the upper bound on sample complexity in Theorem 1 is modified in the presence of loss (see SM S2 C for the proof):
Theorem 3.
[…]
Thus, when $|\beta|^{2}\leq\kappa n$ with a constant $\kappa>0$, $T_{b}=1-O(1/n)$, $T_{a}=1-O(1/n)$ and $r=\Omega(\log n)$, the sample complexity becomes $N=O(\epsilon^{-2}\log\delta^{-1})$ as in the lossless case. For practically relevant squeezing and including loss prior to Bell measurement, we compare the sample complexity for the lossy TMSV+BM protocol and the lossless entanglement-free lower bound in Fig. 3, finding a significant entanglement-enabled advantage in realistic experimental settings. Specifically, for reasonable parameter choices such as squeezing parameter $r=1$, loss rate $10\%$, and $\kappa=O(1)$, we can achieve a factor of $10^{4}~{}(10^{8})$ advantage for around $n=30~{}(60)$ modes. Although the $10^{9}$ number of samples required to achieve the advantage seems large, the state-of-the-art quantum optics experiments (e.g., Refs. [36, 37]) can attain such number of samples in a reasonable time with high sampling rate up to 160 GHz.
Discussion.— We proved that schemes that exploit entanglement with an ancillary quantum memory can learn $n$-mode bosonic random displacement channels with exponentially fewer samples compared to entanglement-free schemes. Our results show that the information-theoretic framework for learning studied in DV quantum systems [17, 19] can be generalized to the CV setting and have powerful implications. We anticipate that these techniques can be applied to other CV learning tasks as well. In addition, our analysis suggests that the separation in sample complexity between entanglement-assisted and entanglement-free protocols may be realized in the near future.
Apart from their theoretical interest, random displacement channels can also be practically relevant in, e.g., modeling noise in bosonic systems. As in the qubit case [38], we expect that noise tailoring methods can transform more general noise models into random displacement channels; therefore efficiently learning random displacement channels can be useful for benchmarking CV quantum systems [39, 40] and for error mitigation.
Displacement estimation is also studied in quantum metrology (see e.g. [41, 42, 43]). A task often considered in metrology is learning an unknown unitary displacement or phase transformation acting independently on each mode [44, 42, 36, 45, 46, 47] whereas the task analyzed in this word is learning an unknown mixture of multimode displacements. Furthermore, while the goal in metrology is typically to learn one or a few parameters, in our case, the parameter space is very large. Therefore, the methodology in the two settings is quite different. Connections between metrology and bosonic channel learning are worthy of further exploration.
Acknowledgements.
We thank Mankei Tsang, Yuxin Wang, Ronald de Wolf, Mingxing Yao, Ming Yuan for insightful discussions. C.O., S.C., Y.W., L.J. acknowledge support from the ARO(W911NF-23-1-0077), ARO MURI (W911NF-21-1-0325), AFOSR MURI (FA9550-19-1-0399, FA9550-21-1-0209), NSF (OMA-1936118, ERC-1941583, OMA-2137642), NTT Research, Packard Foundation (2020-71479). J.P. acknowledges support from the U.S. Department of Energy Office of Science, Office of Advanced Scientific Computing Research (DE-NA0003525, DE-SC0020290), the U.S. Department of Energy, Office of Science, National Quantum Information Science Research Centers, Quantum Systems Accelerator, and the National Science Foundation (PHY-1733907). The Institute for Quantum Information and Matter is an NSF Physics Frontiers Center. S.Z. acknowledges funding provided by the Institute for Quantum Information and Matter and Perimeter Institute for Theoretical Physics, a research institute supported in part by the Government of Canada through the Department of Innovation, Science and Economic Development Canada and by the Province of Ontario through the Ministry of Colleges and Universities. J.A.H.N, Z.L., J.S.N and U.L.A acknowledge support from DNRF (bigQ, DNRF142), IFD (photoQ) and EU (CLUSTEC, ClusterQ ERC-101055224, GTGBS MC-101106833).